# Learning in Structured Stackelberg Games

### v1

@@ A.2 Using multiclass SOA as a black-box algorithm
4:     Commit to $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
@@
Here SOA refers to the online multiclass learning algorithm proposed by [17].
A.3 Online Learning in the Agnostic Setting
A.3.1 Upper bound
@@
See 3.15
@@
First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$. In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate:
@@
From the regret analysis of the Hedge algorithm (see [14] Chapter 2), we know that:
@@
Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$. For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\}$.
For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$. Expert $E_{S,g}$ immitates an instance of the SSOA algorithm, where SSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$. We show the implementation of each expert in Algorithm 4 below.
@@
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
@@
8:     Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Set $l_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
@@
Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type. So, the number of experts that can be constructed in this way is:
$\sum_{i=0}^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}.$
So using the above regret upper bound with this set of experts we get that Algorithm 3 achieves utility at least:
$\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)},$
where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$. It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$. Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that SSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
A.3.2 Lower bound
See 3.16
@@ A.3.2 Lower bound
We let $d=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$. Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$. By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal. In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 00, whereas playing the optimal strategy against the correct follower type gives utility $1$ (for an explicit construction see proof of Theorem 3.8). Then, we have that for every $i$:
@@
We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement. From the SL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$. For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
@@
For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$. So equivalently, we want to lower bound:
@@
Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then:
@@
Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that:
@@
where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [14]). Now it suffices to sum over all $i\in[d]$ to get:
@@
Appendix B Details and Proofs from Section 4
See 4.7

### v2

@@ B.1.3 Using multiclass SOA as a black-box algorithm
4:  Commit to $\mathbf{x}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
@@
Lemma B.6.
There exists a hypothesis class $\pazocal{H}$ and a Stackelberg game $\pazocal{G}$ such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{Ldim}(\pazocal{H})$. Further, SOA achieves strictly larger regret than MSSOA, which in turn achieves strictly larger regret than SSOA.
Proof.
We revisit the game presented in the proof of Theorem˜3.11. We have shown that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$ and $\textnormal{Ldim}(\pazocal{H})=2$. We will now show that $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})=1$. We consider any possible MSL tree. First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered MSL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$). So it suffices to consider contexts 111 and 222. In Figure˜4 we show the two $\pazocal{H}$-shattered MSL tree constructions with $z=1$ and $z=2$ as the root nodes respectively. Notice that the maximal depth of a full binary tree is 111, as desired.
We now prove the regret values. First, since $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ tightly characterizes the regret of SSOA, we know that the regret of SSOA is 3/43/43 / 4. Also, from Theorem˜3.11 we immediately get that the cumulative regret of SOA equals 2 when the sequence of contexts is 111, followed by 222. We now proceed with the regret of MSSOA for the same sequence.
  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$. Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth 111, while the second branch cannot lead to a full binary tree. So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no instantaneous regret).
  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree. SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by MSSOA (thereby incurring instantaneous regret 1).
Figure 4: $\pazocal{H}$-shattered MSL-Tree with (a) $z=1$ and (b) $z=2$ as the root node.
Theorem B.7.
For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes. Moreover, this algorithm coincides with MSSOA.
Proof.
Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
Notice that $\zeta$ is a bijection given that each follower type is distinct (cf. Remark˜2.1). Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$. This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (cf. Theorem 1 in [26]). Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
Consider any MSL-Tree for $\pazocal{H}$. Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well. Also, for any root-to-leaf path on this tree that is consistent with a function $h\in\pazocal{H}$, there exists $\pi_{h}\in\Pi$ that is consistent with it as well, by definition of $\pi_{h}$. The converse is also true: in any Littlestone tree for $\Pi$ consider any root-to-leaf path, that is consistent with some $\pi_{h}\in\Pi$. Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it. Therefore, a tree is MSL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$. This holds for all trees, so it will hold for the maximal-depth trees as well.
B.1.4 No-regret learning in the agnostic setting
Definition B.8 ((Agnostic Setting) Contextual Stackelberg Regret).
where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
@@ B.1.4 No-regret learning in the agnostic setting
Combining our MSSOA algorithm with the above results and tools from multiclass prediction [22], we get the following regret bound.
Theorem B.9.
Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$. In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
@@
First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$. In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate
@@
From the regret analysis of the Hedge algorithm (see [19] Chapter 2), we know that
@@
Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$. For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})\}$.
For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$. Expert $E_{S,g}$ immitates an instance of the MSSOA algorithm, where MSSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$. We show the implementation of each expert in Algorithm˜7 below.
@@
5:  Let $\mathbf{x}^{(h,t)}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
@@
8:  Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:  Set $l_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
@@
Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type. So, the number of experts that can be constructed in this way is $\sum_{i=0}^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}$. So using the above regret upper bound with this set of experts we get that Algorithm˜6 achieves utility at least $\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)}$, where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$. It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$. Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that MSSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
Theorem B.10.
Let $\pazocal{H}$ be any hypothesis class. There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite MSL dimension $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
@@
We let $d=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$. For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$. Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$. By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal. In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 0, whereas playing the optimal strategy against the correct follower type gives utility 111 (for an explicit construction see proof of Theorem˜B.3). Then, we have that for every $i$
@@
We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement. From the MSL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$. For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
@@
For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$. So equivalently, we want to lower bound
@@
Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then
@@
Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that
@@
where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [19]). Now it suffices to sum over all $i\in[d]$ to get
@@
B.2 Distributional Learning
We now give similar results in the distributional setting, in which the leader has access to a batch of training data. We start with some key definitions.
Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times[K]$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as
and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times[K])^{m}$ as
The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times[K])^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to[K]$). We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as
Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as
$\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be
Remark B.11.
For the remainder of this section, we will assume that for any context, $\mathbf{z}$, and follower type $f^{(i)}$ there exists a fixed and known tie-breaking rule over the leader’s optimal strategy against $f^{(i)}$ given $\mathbf{z}$.
Definition B.12 (Natarajan Dimension).
Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to[K]$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have
B.2.1 Vanilla ERM learner
We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$. Formally, we say
Proposition B.13.
Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy
where $h_{ERM}$ satisfies Equation˜3. Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,